OpenStudy (anonymous):

A dozen eggs and one pound of butter cost 1.44. Nine dozen eggs cost as much as 7 pounds of butter. Find the cost of a dozen eggs

OpenStudy (anonymous):

63 cents

OpenStudy (anonymous):

E+B=144 E=144-B 9E=7B 9(144-B=7B 1296-9B=7B 1296=7B+9B 1296=16B B=81 E+81=144 E=144-81 E=63